OpenStudy (anonymous):

What is the value of b? You may assume that the ray is tangent to the circle.

OpenStudy (anonymous):

help

hero (hero):

So how many angles have you found so far?

OpenStudy (anonymous):

3

OpenStudy (anonymous):

@Hero

hero (hero):

What are they? Tell me what you have found so far.

OpenStudy (anonymous):

A B C

hero (hero):

What are the angle measures you have found so far?

OpenStudy (anonymous):

54 86

hero (hero):

That's information that's already given. Is English your first language?

OpenStudy (anonymous):

no

hero (hero):

That's what I figured.

OpenStudy (anonymous):

your rudde why

hero (hero):

angle C = 86/2 angle A = 54/2 Angle B = 90 - angle A

hero (hero):

I promise you I'm not being rude.

OpenStudy (anonymous):

63?

hero (hero):

Good job

OpenStudy (anonymous):

can you help me with another one

hero (hero):

Yeah, post in in another question.